In a certain city the local water utility company charges each customer a flat fee of $4.73 per month plus $2.18 per 1,000 gallons of water used during the month. If the company has 65,000 customers, find the following where x is the total number of gallons used measured in thousands.

A)Find the monthly revenue r for water utility company in terms of x. (use the proper variables and enter as an equation)

B) Find the monthly revene when the average monthly usage per customer is 3,800 gallons of water. Monthly revenue =

Answer by Theo(13342)

from each customer, the local water utility charges:

4.73 + 2.18 * x

that's a flat fee of 4.73 per month plus a variable fee of 2.18 * the number of gallons used measured in thousands.

since the company has 65,000 customers, then their total revenue equation for the month is:

r = 65,000 * (4.73 + 2.18 * x)

r is the total revenue that the company makes each month.
4.73 is the flat fee per customer.
x is the average number of gallons used by each customer divided by 1000.

if the average monthly use of water by each customer during the month is 3800 gallons of water, then the equation would become:

r = 65,000 * (4.73 + 2.18 * 3.8) = 845,910 dollars.

there are 65,000 customers.
each customer pays a flat fee of 4.73.
in addition, each customer pays 2.18 for every thousand gallons of water consumed.
the average number of gallons consumed by each customer is 3800.
3800 / 1000 is equal to 3.8 thousand gallons of water consumed.
the cost per thousand gallons is 2.18, therefore the average cost for each customer is 2.18 * 3.8 = 8.284 dollars for the water consumed.
each customer's average total cost is 4.73 + 8.284 = 13.014
65000 * 13.014 = 845,910.